Cantieri Truck Manufacturing Company (CTMC) ’s balance sheet (Capital structure Section) is given below as of today, April 2, 2018. Ms. Laura Prati, finance manager of the company, has decided to use book value approach to estimate company’s WACC.

Items

Capital Structure

Long Term Debt (par value $1,000)

$ 54,660,000

Preferred Stock

   10,025,000

Common Stock

   99,940,000

Total

164,625,000

CTMC has outstanding bonds with an annual 8 percent coupon. The bonds have a par value of $1,000, will mature in 11 years, and currently are selling at $865. The preferred stock of the company pays $2.50 preferred dividend per share per year and is currently trading at $21.65 per share at the Sassari Stock Exchange (SSE).

Ms. Prati has collected data to indicate the risk-free rate to be 5 percent and expected return on market portfolio to be 12.30%. CTMC beta is estimated to be 1.25.

CTMC stocks are traded at $50.50 per share and the company recently paid dividend $4.00 per share. Financial analysts believe CTMC will grow at a constant rate of 6% in future years.

CTMC’s current policy is to add a risk premium of 4.25% to the company’s own before tax cost of debt, when bond-yield-plus-risk-premium approach is used to estimate company’s own cost of equity. CTMC tax rate is 30 percent.

1) Use book value approach to estimate company’s WACC. Show work.
Answer #1

use rate formuale in excel to find the ytm of the bond

Book value Weights Cost of indiv WACC Debt Preferred stock Common stock Totak 54,660,000| 33.20% 6.09% 99,940,000| 60.71% 164

=rate(nper,pmt,pv,fv,type)
=rate(11,(54660000*8%),(-865*(54660000/1000)),54660000,0)
=10.09%
After tax cost of debt=10.09%*(1-30%)=7.06%
Cost of preferred stock= preference dividend/price of stock
=2.5/21.65
=11.55%
Since cost of equity given 3 methods we take average of it
1st method CAPM
=Risk free+(beta*(market return-riskfree))
=5%+(1.25*(12.3%-5%))
=14.13%
2nd method: dividend growth
stock price=D*(1+g)/(k-g)
50.5=4*(1+6%)/(k-6%)
k=6.84%
3rd method:
=10.09%+4.25%=14.34%
Average Cost of equity=(14.13%+6.84%+14.34%)/3=11.77%
WACC=(7.06%*